After driving around and undergoing a displacement of 27.9 km and N 16.5 km, a driver finds herself at position r2x--39.3 km and rzy 16.7 km What were the components ofthe drivers initial position vector, rix and ry, in kilometers? Number km 1x Number km
Answer #1

x-component of driver's initial position is r1x = -39.3+27.9 = -11.4 km


y-component of driver's initial position is r1y = 16.7-16.5 = 0.2 km